Judge Jackie LLC expects its lawyers to work a total of 21,000 direct labor hours during the upcoming year and are paid $525,000. Estimated total indirect costs are $252,000. The indirect cost allocation rate is based on direct labor hours.

What is the total cost of Job 628 that used 12 hours of labor time?

Round to the nearest whole number. Do not include dollar signs or commas in your answer.
Answer #1

answer:444

total cost=direct labour cost of 12 hours+indirect cost for 12 hours

direct labour cost per hour=525000/21000=25 per hour

direct labour cost for 12 hours=25*12=300

indirect cost rate=252000/21000=12 per hour

indirect cost for 12 hours=12*12=144

total cost=300+144=444